OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y = 1/4x + 2. y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x

OpenStudy (anonymous):

Hey there :) .. I believe I can answer your question .. But I just don't see it clearly.... In the question. You have '4x+2' there next to 'y=1' .... Is that all part of one equation or what?

OpenStudy (anonymous):

K so with a prependicular line, to find the slope we use the negative reciprocal. What that means is for instance we have 5, our negative reciprocal would be -1/5. We flipped the number around and flipped the sign. For this equation, the slope is -4 and we put the equation into point-slope formula. y-y1=m(x-x1) y+2=-4(x+1) y+2=-4x-4 [Subtract 2 on both sides] y=-4x-6, which is your answer B